# Birch SGD: A Tree Graph Framework for Local and Asynchronous SGD Methods

### v2

@@ 3 Existing and New Algorithms: Summary and Comparison
3. Communication complexity. Communication delays are a major bottleneck in real-world distributed systems. Thus, minimizing communication and synchronization overhead is crucial. Ringmaster ASGD is the least efficient in this regard, requiring frequent communication and lacking AllReduce support due to its highly asynchronous design (see Alg. 7). In a simple model where sending one stochastic gradient takes $\tau$ seconds and all workers have identical speed $h_{i}=h$, the time complexity of Ringmaster ASGD is $\Omega\left(\tau\left(\nicefrac{{L\Delta}}{{\varepsilon}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}}\right)+h\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}\right)\right).$ In contrast, Rennala SGD achieves $\mathcal{O}\left(\tau\times\nicefrac{{L\Delta}}{{\varepsilon}}+h\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}\right)\right),$ which is preferable when $\tau$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ are large (see Section G).
@@
Figure 4: An example of a Local SGD computation tree with $B=4$ and 2 workers, each performing local steps over 2 global steps. In first round, they calculate $M_{1}=2$ and $M_{2}=2$ local steps. In the second round, they calculate $M_{1}=1$ and $M_{2}=3$ local steps. Note that the maximum distances $\textnormal{dist}(x^{3},z_{1}^{0,1})$ and $\textnormal{dist}(x^{7},z_{2}^{1,2})$, when applying $\nabla f(z_{1}^{0,1};\eta^{0,1}_{1})$ to $x^{3}$ and $\nabla f(z_{2}^{1,2};\eta_{2}^{1,2})$ to $x^{7}$, are equal to $B-1=\sum_{i=1}^{n}M_{i}-1=3$. Notice that each stochastic gradient is used $2$ times in the tree.
We consider the classical Local SGD strategy, where each worker $i\in[n]$ performs $M_{i}$ local steps, after which the server aggregates the results (see Alg. 5). Unlike most previous approaches, however, the number of local steps ${M_{i}}$ may vary across workers. Moreover, the server waits for a specific condition before aggregating: $\sum_{i=1}^{n}M_{i}=B.$ This strategy is adaptive to fluctuations in the number of local steps performed by individual workers, as the server ensures the total number of steps across all workers reaches the target sum $\sum_{i=1}^{n}M_{i}=B,$ where $B$ is a hyperparameter. An example of the computation tree is shown in Figure 4.
In Section E.3, we establish the iteration rate of Local SGD as $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$. This result follows directly from Theorem 2.4 via a simple geometric argument. In fact, looking at Figure 4 reveals that all the conditions of Theorem 2.4 are satisfied. The only minor difficulty is to show that $R:=\sup_{k\geq 0}\textnormal{dist}(x^{k},z^{k})\leq B-1,$ which is guaranteed by the condition $\sum_{i=1}^{n}M_{i}=B.$
What is novel in our version of Local SGD is the stopping condition and the choice of $B$, which together ensure the optimality of Local SGD under the $h_{i}$-fixed computation model (see Thm. F.6).
@@
Dual-Process SGD (Section E.11). We took a step further and developed a new local method inspired by Local SGD and Async-Local SGD. It is the first local method to achieve the optimal time complexity in the heterogeneous time setting, where workers have varying computation and communication times (see Section H). However, unlike Local SGD, it is not AllReduce-friendly.
Cycle SGD (Section E.5). To mitigate this issue, we propose a new method called Cycle SGD. Similar to Local SGD, each worker performs local steps. However, once the workers finish computing the initial stochastic gradients $\{\nabla f(z^{0}_{i};\eta^{0}_{i})\}$, only the first group of $s$ workers sends their gradients to the server, where $s$ is a hyperparameter. The server then aggregates these gradients and performs the update $w^{1}=w^{0}-\gamma\sum_{i=1}^{s}\nabla f(z^{0}_{i};\eta^{0}_{i}).$ Meanwhile, the first $s$ workers begin computing their local steps starting from $w^{1}$, while the remaining workers continue their current local computations. Subsequently, the second group of $s$ workers sends their locally computed vectors, and this process continues in a circular manner. A computation tree presented in Figure 16. In Section E.5, we provide a formal descriptions and, using Theorem 2.4, prove that Cycle SGD converges after $\mathcal{O}\left(\nicefrac{{n^{2}L\Delta}}{{s\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$ iterations. The peak bandwidth of Cycle SGD is $\mathcal{O}\left(s\right)$ with $s=\min\left\{\max\{\lceil\nicefrac{{n^{2}\varepsilon}}{{\sigma^{2}}}\rceil,1\},n\right\},$ which is much smaller than $\Theta\left(n\right)$ when $\nicefrac{{\sigma^{2}}}{{\varepsilon}}\geq n.$
5. Optimization with clusters. Consider a scenario with two or more clusters of workers, where intra-cluster communication (e.g., InfiniBand) is fast, but inter-cluster communication (e.g., Ethernet) is slow.
@@
6. Flexible synchronization. We noticed that in all previous methods, the workers are synchronized in a predefined manner or rule. We want to add more flexibility to the synchronization process.
Meta Local SGD (Section E.10). Our idea is that the server (or the workers themselves, in a decentralized setup) can select any subset of workers based on any strategy (e.g., randomly or according to current communication speeds), gather their computed stochastic gradients, update the global model, and ask these workers to continue performing local steps from the new point. However, such “anarchic synchronization” can result in a computation tree with a large $R$ if the selected strategy is not chosen carefully. To ensure that $R$ is bounded, in our meta-algorithm (Algorithm 16), we track the current distances $\{d_{i}\}$ to the head of the main branch and the local steps $\{M_{i}\}$ performed by each worker. Then, by tracking the value $d_{i}+\sum_{i=1}^{n}M_{i}$ for all $i\in[n]$ and comparing it to a parameter $B$, we compulsorily synchronize (Hard Sync) all workers for which $d_{i}+\sum_{i=1}^{n}M_{i}=B$. This way, we can ensure that the maximum distance $R$ is bounded by $B$, and the iteration rate of this method is $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$. See details in Section E.10.
@@ 4 Summary of experimental results
In Section C, we provide a detailed comparison of methods on logistic regression, image classification with ResNet18 (He et al.,, 2016), and next-token prediction with GPT2 (Radford et al.,, 2019). When communication times are negligible (Fig. 7), as expected from Table 1 and the previous discussion, Ringmaster ASGD and Async-Local SGD converge faster on the logistic regression problem. However, when communication times are large (Fig. 7), Ringmaster ASGD becomes less practical due to its frequent updates. Synchronized SGD exhibits the worst performance across all setups. Rennala SGD and Local SGD are more stable, while Async-Local SGD performs well due to its effective balance between frequent updates and local steps.
@@
5 Insights, Guidelines, and Conclusion
All proposed methods share the same iteration rate of $\mathcal{O}\left(\nicefrac{{(R+1)L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right),$ where $R$ is controlled by a method-specific hyperparameter and, at the same time, $R$ is the largest tree distance between $x^{k}$ and $z^{k}$. For Rennala SGD, $R=B-1,$ where $B$ denotes the batch size; for Ringmaster ASGD, $R=B-1,$ where $B$ is the delay threshold; for Local SGD, $R=B-1,$ where $B$ corresponds to the number of local steps; for Cycle SGD, $R=\nicefrac{{n^{2}}}{{s}},$ where $s$ is the group size, etc. In all these methods, $R$ can be controlled, and to achieve the best possible computational and communication guarantees, one should always choose $R=\Theta\left(\nicefrac{{\sigma^{2}}}{{\varepsilon}}\right)$ (see Sections G and F). We believe this is a fundamental principle underlying all parallel optimization methods, and it should be considered a guiding rule when developing new algorithms. This choice is also theoretically justified: by taking $R=\Theta\left(\nicefrac{{\sigma^{2}}}{{\varepsilon}}\right)$, the iteration rate does not change asymptotically: $\mathcal{O}\left(\nicefrac{{(R+1)L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)=\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$. Larger values of $R$ allow the methods to be more “parallel-friendly”. For instance, a large $R$ enables Ringmaster ASGD to consider stochastic gradients with larger delays, while a large $R$ in Local SGD allows the method to run more local steps. However, taking $R>\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ results in a worse iteration rate, suggesting that the corresponding method operates in an overly “anarchic” asynchronous regime, which may lead to performance degradation. Geometrically, the theory suggests that, to achieve good performance, the tree distance between $x^{k}$ and $z^{k}$ in Figure 3 should not exceed $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$.
We hope these important observations will support the future development and analysis of asynchronous optimization methods. Building on these insights, we designed at least eight new methods using our proposed Birch SGD framework and the main result, Theorem 2.4. By reducing the analysis and design of these methods to computation trees, our entire development becomes purely graph-geometrical, offering a new and simpler view on asynchronous optimization methods.
@@ A.1 Discussion of the computational time complexities
To compare parallel and asynchronous methods, Mishchenko et al., (2022) proposed using the $h_{i}$-fixed computation model. The idea is to assume that worker $i$ requires at most $h_{i}$ seconds to calculate one stochastic gradient for all $i\in[n]:=\{1,\dots,n\}$ (without loss of generality, $h_{1}\leq h_{2}\leq\dots\leq h_{n}$). The authors considered Synchronized SGD, an iterative process defined as $w^{k+1}=w^{k}-\frac{\gamma}{n}\sum_{i=1}^{n}\nabla f(w^{k};\eta^{k}_{i}),$ where each worker calculates one stochastic gradient, synchronize, and a parameter server aggregates them to update the iterate [^4] Alternatively, there is no physical parameter server, and all workers perform an Allreduce.. Using the $h_{i}$-fixed computation model, it can be easily shown that Synchronized SGD converges after
@@
Mishchenko et al., (2022); Koloskova et al., (2022) provided new analyses of Asynchronous SGD (see Algorithm 2) and Cohen et al., (2021) developed Picky SGD to show that this time complexity can be improved to
@@
which is achieved by the Rennala SGD method (Tyurin and Richtárik,, 2023). Moreover, Tyurin and Richtárik, (2023) proved a matching lower bound demonstrating that both this complexity and Rennala SGD are optimal. Recently, Maranjyan et al., (2025) developed a new optimal Ringmaster ASGD method, which is essentially Asynchronous SGD with a key modification. Additionally, under the universal computation model, Tyurin, (2025); Maranjyan et al., (2025) showed that both Rennala SGD and Ringmaster ASGD remain optimal even when computation times are arbitrary, time-varying, and random.
@@ A.2 More related work
Our focus is on the homogeneous setting, where all workers have access to the same data distribution or dataset. The heterogeneous data setting is equally important, especially in federated learning (FL) (Konečný et al.,, 2016) due to privacy constraints. In this context, many other methods have been proposed, including Asynchronous SGD (Mishchenko et al.,, 2022; Koloskova et al.,, 2022), Asgrad (Islamov et al.,, 2024), PIAG (Wu et al.,, 2022), and Malenia SGD (Tyurin and Richtárik,, 2023). Notably, Tyurin and Richtárik, (2023); Tyurin, (2025) showed that Malenia SGD is optimal under both the fixed and universal computation models, without requiring assumptions of bounded gradients or gradients dissimilarity.
In the homogeneous setting, numerous other works have studied asynchronous SGD methods, including (Lian et al.,, 2015; Feyzmahdavian et al.,, 2016; Stich and Karimireddy,, 2020; Sra et al.,, 2016). However, these methods typically require the assumption that the delays in the indices of stochastic gradients are bounded (on average in (Sra et al.,, 2016)). As a result, their theoretical guarantees in terms of computational time complexity are weaker than those in (Cohen et al.,, 2021; Koloskova et al.,, 2022; Mishchenko et al.,, 2022; Tyurin and Richtárik,, 2023; Maranjyan et al.,, 2025), which do not rely on such assumptions.

### v3

@@ 3 Existing and New Algorithms: Summary and Comparison
3. Communication complexity. Communication delays are a major bottleneck in real-world distributed systems. Thus, minimizing communication and synchronization overhead is crucial. Ringmaster ASGD is the least efficient in this regard, requiring frequent communication and lacking AllReduce support due to its asynchronous design. In a simple model where sending one stochastic gradient takes $\tau$ seconds and all workers have identical speed $h_{i}=h$, the time complexity of Ringmaster ASGD is $\Omega\left(\tau\left(\nicefrac{{L\Delta}}{{\varepsilon}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}}\right)+h\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}\right)\right).$ In contrast, Rennala SGD achieves $\mathcal{O}\left(\tau\nicefrac{{L\Delta}}{{\varepsilon}}+h\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{n\varepsilon^{2}}}\right)\right)$ (see Sec. G), which is better when $\tau$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ are large.
@@
$\overset{\shortstack{\text{Sync. via server}\\ \text{or {AllReduce}}}}{\Rightarrow}$
$\overset{\shortstack{\text{After two global steps}}}{\Rightarrow}$
Figure 4: An evolution of a Local SGD computation tree with $B=4$ and 2 workers, each performing local steps over 2 global steps. In the first round, they compute $M_{1}=2$ and $M_{2}=2$ local steps (first figure), after which they synchronize (second figure). In the second round, they compute $M_{1}=1$ and $M_{2}=3$ local steps and synchronize again (third figure). Note that the maximum distances $\textnormal{dist}(x^{3},z_{1}^{0,1})$ and $\textnormal{dist}(x^{7},z_{2}^{1,2})$, when applying $\nabla f(z_{1}^{0,1};\eta^{0,1}_{1})$ to $x^{3}$ and $\nabla f(z_{2}^{1,2};\eta_{2}^{1,2})$ to $x^{7}$, are equal to $B-1=\sum_{i=1}^{n}M_{i}-1=3$. Notice that each stochastic gradient is used 22 times in the tree.
We consider the classical Local SGD strategy, where each worker $i\in[n]$ performs $M_{i}$ local steps, after which the server aggregates the results (see Alg. 5). Unlike most previous approaches, however, the number of local steps $M_{i}$} may vary across workers. Moreover, the server waits for a specific condition before aggregating: $\sum_{i=1}^{n}M_{i}=B.$ This strategy is adaptive to fluctuations in the number of local steps performed by individual workers, as the server ensures the total number of steps across all workers reaches the target sum $\sum_{i=1}^{n}M_{i}=B,$ where $B$ is a hyperparameter. An example of the computation tree is shown in Figure 4. In Section E.3, we establish the iteration rate of Local SGD as $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$. This result follows directly from Theorem 2.4 via a simple geometric argument. In fact, looking at Figure 4 reveals that all the conditions of Theorem 2.4 are satisfied. The only minor difficulty is to show that $R:=\sup_{k\geq 0}\textnormal{dist}(x^{k},z^{k})\leq B-1,$ which is guaranteed by the condition $\sum_{i=1}^{n}M_{i}=B.$
What is novel in our version of Local SGD is that it achieves better theoretical guarantees within the family of Local SGD approaches (see Section H). Moreover, our stopping condition and the choice of $B$ together ensure its optimality under the $h_{i}$-fixed computation model (see Theorem F.6).
@@
Dual-Process SGD (Section E.11). We took a step further and developed a new local method inspired by Local SGD and Async-Local SGD. It is the first local method to achieve the optimal time complexity in the distributed setting, where workers have varying computation and communication times (see Section I). However, unlike Local SGD, it is not AllReduce-friendly.
Cycle SGD (Section E.5). To mitigate this issue, we propose a new method called Cycle SGD. Similar to Local SGD, each worker performs local steps. However, once the workers finish computing the initial stochastic gradients $\{\nabla f(z^{0}_{i};\eta^{0}_{i})\}$, only the first group of $s$ workers sends their gradients to the server, where $s$ is a hyperparameter. The server then aggregates these gradients and performs the update $w^{1}=w^{0}-\gamma\sum_{i=1}^{s}\nabla f(z^{0}_{i};\eta^{0}_{i}).$ Meanwhile, the first $s$ workers begin computing their local steps starting from $w^{1}$, while the remaining workers continue their current local computations. Next, the second group of $s$ workers sends their locally computed vectors, and this process continues in a circular manner. A computation tree presented in Figure 18. The peak bandwidth of Cycle SGD is $\mathcal{O}\left(s\right)$ with $s=\min\left\{\max\{\lceil\nicefrac{{n^{2}\varepsilon}}{{\sigma^{2}}}\rceil,1\},n\right\},$ which is smaller than $\Theta\left(n\right)$ when $\nicefrac{{\sigma^{2}}}{{\varepsilon}}\geq n.$
5. Optimization with clusters. Consider a setup with many clusters of workers, where intra-cluster communication (e.g., InfiniBand) is fast and inter-cluster communication (e.g., Ethernet) is slow.
@@
6. Flexible synchronization and Meta Local SGD (Section E.10). We noticed that in all previous methods, the workers are synchronized in a predefined manner or rule. We want to add more flexibility to the synchronization process. Our idea is that the server (or the workers themselves, in a decentralized setup) can select any subset of workers based on any strategy (e.g., randomly or according to current communication speeds), gather their computed stochastic gradients, update the global model, and ask these workers to continue performing local steps from the new point. However, such “anarchic synchronization” can result in a computation tree with a large $R$ if the selected strategy is not chosen carefully. To ensure that $R$ is bounded, in our meta-algorithm (Algorithm 16), we track the current distances $\{d_{i}\}$ to the head of the main branch and the local steps $\{M_{i}\}$ performed by each worker. Then, by tracking the value $d_{i}+\sum_{i=1}^{n}M_{i}$ for all $i\in[n]$ and comparing it to a parameter $B$, we compulsorily synchronize (Hard Sync) all workers for which $d_{i}+\sum_{i=1}^{n}M_{i}=B$. This way, we can ensure that $R$ is bounded by $B$, and the iteration rate of this method is $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$.
4 Insights and Guidelines
All proposed methods share the same iteration rate of $\mathcal{O}\left(\nicefrac{{(R+1)L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right),$ where $R$ is controlled by a method-specific hyperparameter and, at the same time, $R$ is the largest tree distance between $x^{k}$ and $z^{k}$. For Rennala SGD, R=B1,R=B-1, where $B$ denotes the batch size; for Ringmaster ASGD, R=B1,R=B-1, where $B$ is the delay threshold; for Local SGD, R=B1,R=B-1, where $B$ corresponds to the number of local steps; for Cycle SGD, $R=\nicefrac{{n^{2}}}{{s}},$ where $s$ is the group size, etc. In all these methods, $R$ can be controlled, and to achieve the best possible computational and communication guarantees, one should always choose $R=\Theta\left(\nicefrac{{\sigma^{2}}}{{\varepsilon}}\right)$ (see Sections G and F). We believe this is a fundamental principle underlying all parallel optimization methods, and it should be considered a guiding rule when developing new algorithms. This choice is also theoretically justified: by taking $R=\Theta\left(\nicefrac{{\sigma^{2}}}{{\varepsilon}}\right)$, the iteration rate does not change asymptotically: $\mathcal{O}\left(\nicefrac{{(R+1)L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)=\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$. Larger values of $R$ allow the methods to be more “parallel-friendly”. For instance, a large $R$ enables Ringmaster ASGD to consider stochastic gradients with larger delays, while a large $R$ in Local SGD allows the method to run more local steps. However, taking $R>\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ results in a worse iteration rate, suggesting that the corresponding method operates in an overly “anarchic” asynchronous regime, which may lead to performance degradation. Geometrically, the theory suggests that, to achieve good performance, the tree distance between $x^{k}$ and $z^{k}$ in Figure 3 should not exceed $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$.
While our primary focus is on training large-scale language and vision models, where the i.i.d. assumption is usually appropriate, we acknowledge that non-i.i.d. scenarios are also important and should be investigated in future work, where it may be necessary to add additional assumptions such as first-order and second-order similarity of the functions (Arjevani and Shamir, 2015; Mishchenko et al., 2022). Moreover, it would be interesting to extend our framework to methods with preconditioning (Kingma and Ba, 2015) and non-Euclidean geometry (Jordan et al., ).
We hope our observations will support the future development and analysis of asynchronous optimization methods. Building on these insights, we designed at least eight new methods using our proposed Birch SGD framework and the main result, Theorem 2.4. By reducing the analysis and design of these methods to computation trees, our entire development becomes purely graph-geometrical, offering a new and simpler view on asynchronous optimization methods.
@@ 5 Summary of experimental results
In Section C, we provide a detailed comparison of methods on logistic regression, image classification with ResNet18 (He et al., 2016), and next-token prediction with GPT2 (Radford et al., 2019). When communication times are negligible (Fig. 7), as expected from Table 1 and the previous discussion, Ringmaster ASGD and Async-Local SGD converge faster on the logistic regression problem. However, when communication times are large (Fig. 7), Ringmaster ASGD becomes less practical due to its frequent updates. Synchronized SGD exhibits the worst performance across all setups. Rennala SGD and Local SGD are more stable, while Async-Local SGD performs well due to its effective balance between frequent updates and local steps.
@@
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement №139-10-2025-033.
@@ A.1 Discussion of the computational time complexities
To compare parallel and asynchronous methods, Mishchenko et al. (2022) proposed using the $h_{i}$-fixed computation model. The idea is to assume that worker $i$ requires at most $h_{i}$ seconds to calculate one stochastic gradient for all $i\in[n]:=\{1,\dots,n\}$ (without loss of generality, $h_{1}\leq h_{2}\leq\dots\leq h_{n}$). The authors considered Synchronized SGD, an iterative process defined as $w^{k+1}=w^{k}-\frac{\gamma}{n}\sum_{i=1}^{n}\nabla f(w^{k};\eta^{k}_{i}),$ where each worker calculates one stochastic gradient, synchronize, and a parameter server aggregates them to update the iterate [^4] Alternatively, there is no physical parameter server, and all workers perform an Allreduce.. Using the $h_{i}$-fixed computation model, it can be easily shown that Synchronized SGD converges after
@@
Mishchenko et al. (2022); Koloskova et al. (2022) provided new analyses of Asynchronous SGD (see Algorithm 2) and Cohen et al. (2021) developed Picky SGD to show that this time complexity can be improved to
@@
which is achieved by the Rennala SGD method (Tyurin and Richtárik, 2023). Moreover, Tyurin and Richtárik (2023) proved a matching lower bound demonstrating that both this complexity and Rennala SGD are optimal. Recently, Maranjyan et al. (2025) developed a new optimal Ringmaster ASGD method, which is essentially Asynchronous SGD with a key modification. Additionally, under the universal computation model, Tyurin (2025); Maranjyan et al. (2025) showed that both Rennala SGD and Ringmaster ASGD remain optimal even when computation times are arbitrary, time-varying, and random.
@@ A.2 More related work
Our focus is on the homogeneous setting, where all workers have access to the same data distribution or dataset. The heterogeneous data setting is equally important, especially in federated learning (FL) (Konečný et al., 2016) due to privacy constraints. In this context, many other methods have been proposed, including Asynchronous SGD (Mishchenko et al., 2022; Koloskova et al., 2022), Asgrad (Islamov et al., 2024), PIAG (Wu et al., 2022), and Malenia SGD (Tyurin and Richtárik, 2023). Notably, Tyurin and Richtárik (2023); Tyurin (2025) showed that Malenia SGD is optimal under both the fixed and universal computation models, without requiring assumptions of bounded gradients or gradients dissimilarity.
In the homogeneous setting, numerous other works have studied asynchronous SGD methods, including (Lian et al., 2015; Feyzmahdavian et al., 2016; Stich and Karimireddy, 2020; Sra et al., 2016). However, these methods typically require the assumption that the delays in the indices of stochastic gradients are bounded (on average in (Sra et al., 2016)). As a result, their theoretical guarantees in terms of computational time complexity are weaker than those in (Cohen et al., 2021; Koloskova et al., 2022; Mishchenko et al., 2022; Tyurin and Richtárik, 2023; Maranjyan et al., 2025), which do not rely on such assumptions.
A.3 Relation to other frameworks
There were several previous approaches to unify SGD methods. Gorbunov et al. (2021) proposed using a parametric assumption to unify the analysis of local methods, Wang et al. (2020) unified FedAvg-like methods, Wang and Joshi (2018) proposed Cooperative SGD to analyze different synchronization mechanisms through mixing matrices, Huang et al. (2022) analyzed a general sample-wise Push–Pull framework in the heterogeneous setting using two-level augmented graphs, and Khaled et al. (2020); Gorbunov et al. (2020) analyzed SGD methods with variance-reduction and compression techniques. These approaches are related to, but not directly comparable with ours, and, to the best of our knowledge, our approach is new and orthogonal.
Another interesting work that analyzes SGD methods is (Even et al., 2024). Their work and ours both use graphs; however, we use graphs in completely different, orthogonal, and unrelated contexts. In their case, nodes represent computers (GPUs, CPUs, servers), and edges represent communication links. In our case, nodes represent points of an algorithm, (directed) edges indicate how one point is calculated from another, and the graphs evolve with every iteration. Similarly, Huang et al. (2022) also use a graph abstraction but with a different meaning for nodes and edges: in their case, nodes correspond to devices and iterates of data samples, edges represent both communication and computation links, and the number of nodes is fixed from the beginning since every node corresponds to one sample or worker. These are different and orthogonal approaches. Another important difference is that they compare methods using iteration complexities, whereas we use time complexities in Table 1, which is a more robust and suitable metric for asynchronous and parallel methods.